Alfio,Juan Carlos,Enrico, Runa

start up ok.

secondary start up ok.

The center stage didn't move correctly due to a problem of the rerorator homing.

Scopriamo che l'homing del rirotatore era stato disabilitato per sbaglio dal file di config.

PROBLEM.
the secondary didn't accept the shape (RETRY) which it was working yesterday,.Load shape; flat_closeloop_200modesKL.sav

ACT-SUBAP matching

Cerchiamo di identificare acts e sottoaperture nella config buona del FLAO01 per poterla riprodurre.

Pattern di attuatori con geometria di Fried nella conf buona da essere identificati con i segnali del W:
fried_acts_up150nm.fits
fried_acts_down150nm.fits

Attuatori mossi :144 164 311 363 570

Acquisisamo le pup xche' le ultime erano vicine al confine dei quadranti.
Pup = 20090918_103144
Centri = 57.5-21-5
Diam = 30.0
Dist = 36.02pix
TTMod = 8V 200Hz
Thr1 = 0.15
Thr2 = 0.15
Sottoap centrale in slopes di sx: 23;58

Stop test, working on in an improvement of the pupil acquire procedure
NEW ALGORITHM TO BUILD THE PUPILS. < 20090918_150050 >.

- BUILD A MASK WITH A DEFINED RADIUS.
- SHIFT THE MASK TO THE PUPIL CENTER MEASURED.
- insert secondary pattern, rotate and move x or y the to center the central oscuration ( defined central pixel)
- DEFINE THE NEW PUPILS USING THE HOT PROCEDURE.
-
PUPIL ACQUISTION PROCEDURE HOT <20090918_152412>.

Back to the test.

MATCHING SECONDARY ACTUATORS -SUBAPERTURES

Pattern di attuatori con geometria di Fried nella conf buona da essere identificati con i segnali del W:
fried_acts_up150nm.fits
fried_acts_down150nm.fits

Centering using the pupil 20090918_150050 of 31 pixel and applying the file doppio_radial_verticale50_coda.fits, adjusting the rotation found 174.0 deg and translation x of the camera lens. Position of the camera lens: x:0.0 , y:10.0.

Checking this new set up acquiring a new pupils: < 20090918_173713 >

the centers of those new pupils will be used for the camera lens loop.

Adjusting the Y direction using the pattern doppio_radial_orizontale50.fits . Moving the camera lens in Y at position +4.

AFTER ADJUSTING THE X AND Y AND ROTATION A NEW PUPILS ARE ACQUIRED: < 20090918_182832 >, PROCEDURE INCOMPLETE........

It has been done a check with the new centers and the differences is zero.

BOARD SET UP SAVED . WFS_REGISTEREDV1

TEST.

Rotation of the rerotator 180 degrees and come back the postions didn't change the positions ( there are not difference in the positions before and after).

ACQUIRE PUPIL TYPE HOT < 20090918_183032>

ACQUIRE INTMAT 20090918_183443

INTMATACQ <tracking#>
Pup = <>
Dist = <>
Modul =<3 l/D at 400+HZ>

ANALYSE INTMAT. 20090918_

looking the modes they are looking noisy.

gENERATE A NEW DISTURB FILE
CmdDisturb _20090918_184025

Modal base: KL
Nmodes:400.
400Hz.
3l/D.
acquire INTMAT. < 20090918_184107>

Nmodes:400
Cmddist_20090918_184025
400Hz.
3L/D

this int mat couldn't be analysed problem with the frame counter.
Analyse INTMAT. < >wrong

acquire INTMAT. < 20090918_184434>
analyse INTmat . < Rec_20090918_184605>

looking the modes they looks noisy, the test keep going.

TEST CLOSE LOOP WITH THE REC.<Rec_20090918_184605>

OPTICAL LOOP DIAGNOSTIC < 20090918_184828 >

ALL SAVED
N FRAMES SAVED 4000 FRAMES.
DistURBANCE: NO

OPTICAL LOOP DIAGNOSTIC < 20090918_185028 >

Dist: On 80pix. oversampling 800
gain: 0.7

ACQUIRE SLOPE NULL WITH A DEFOCUS.
the close loop didn't go

-- JuanCarlosGuerra - 2009-09-18
Topic revision: r6 - 24 Sep 2009, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback